1. Calculate the equilibrium level of income if C=R100 million +0,8Y and Ī =R125 million .
1
Expert's answer
2019-08-07T09:19:47-0400

The equilibrium income level is determined by the formula, according to the source data:

Y = C + I

Y = 100 + 0.8Y + 125

(1-0.8) Y = (100 + 125)

0.2Y = 225

Y = 225 / 0.2

Y = 1125